OpenStudy (anonymous):

integrate sin^43x dx

OpenStudy (anonymous):

\[\int\limits \sin^{43}(x) dx= - \cos^{43}(x)\] honestly, this is a weird question lol, so I'm not sure? ._.

OpenStudy (anonymous):

+ C ofcourse

OpenStudy (anonymous):

sstar, i don't know the answer but i'm almost sure that's not correct ><

OpenStudy (anonymous):

LOL >_< I know, it looks weird

OpenStudy (anonymous):

hmm, is it possible to let u = 43? I've never solved such thing lol

OpenStudy (anonymous):

meh... no...this method is funny lol

OpenStudy (anonymous):

sstarica the question is\[\int\limits \sin ^{4}3x dx\]

OpenStudy (anonymous):

no, you have to substitute (cos x)^2 = 1 - (sin x)^2

OpenStudy (anonymous):

oh that makes everything a lot better.... oneprince USE PARENTHESIS!

OpenStudy (anonymous):

ok chia

OpenStudy (anonymous):

no wait a min, isn't the sin to the power of 43?

OpenStudy (anonymous):

chia.. solve it

OpenStudy (anonymous):

was it to the power of 43? or to the power of 2? lol

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

it's sin(3x)^4

OpenStudy (anonymous):

that's the question?

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

ugh...sub 3x for u and then use the half angle formula to get cosine and sine together, integrate

OpenStudy (anonymous):

sstar are you writing it out? caz i don't really want to...

OpenStudy (anonymous):

>_< nvm, proceed chia, I misunderstood the question lol

OpenStudy (anonymous):

oneprince, did you read what i wrote? is that enough?

OpenStudy (anonymous):

AMISTRE! you can do this! ^_^

OpenStudy (amistre64):

Howdy :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

my brain is defunctioning, so I don't think I can write it down even though I took it this semester and it looks SIMPLE! would you amistre? ^_^

OpenStudy (amistre64):

[S] sin^4(3x) dx this is missing something lol this came from something similar to -cos^5(3x) so what do we get when we derive that? and itll tell us what to add...

OpenStudy (anonymous):

it's not really missing something, you can let u = 3x and solve normally like chia said

OpenStudy (anonymous):

if only I remember lol >_<

OpenStudy (amistre64):

Dx(-cos^5(3x)) = 5(3) sin^4(3x) which would be easy to solve right? So we need a "15" in front of your integral to convert it

OpenStudy (amistre64):

lets multiply your original problem by "1"; or rather a convenient for of "1" such as 15/15..

OpenStudy (anonymous):

where did you get cos^5?

OpenStudy (amistre64):

then we can pull out the bottom "15" and integrate the rest up like normal

OpenStudy (amistre64):

sin^4 comes from cos^5; or at least a version of it right?

OpenStudy (anonymous):

you can do this : \[\int\limits (1-\cos^2(3x))^2 dx = \int\limits(1-2\cos^2u+\cos^2u)du\] and simply solve, right?

OpenStudy (anonymous):

where u = 3x ? :)

OpenStudy (anonymous):

right? .-.

OpenStudy (anonymous):

isn't it simpler that way, lol?

OpenStudy (amistre64):

\[\int\limits_{} \frac{5(3)}{5(3)} \sin^4 (3x) dx\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

amistre, i think you're doing it wrong.. and i feel bad caz i'm not actually doing it...

OpenStudy (amistre64):

\[\frac{1}{15} \int\limits_{} 5(3) \sin^4 (3x) dx\]

OpenStudy (amistre64):

-cos^5 (3x)/15 + C

OpenStudy (anonymous):

use the half angle formula to integrate sin or cos with a power you MUST have sin * cos

OpenStudy (anonymous):

no, amistre, that's definitely wrong T-T

OpenStudy (anonymous):

Why don't you make a substitution of \[u=3x \rightarrow dx = \frac{du}{3}\]so that then\[I=\frac{1}{3}\int\limits_{}{}\sin^4 u du\]and now use a reduction formula?

OpenStudy (amistre64):

derive it back again, its right.... it should be right lol

OpenStudy (anonymous):

using the way I cut it down, you'll get : \[= \int\limits1 du - 2\int\limits \cos^2u du + \int\limits \cos^2u du\] then you'll use for cos^2 u = 1/2(1+cos2u)

OpenStudy (anonymous):

star is right

OpenStudy (anonymous):

^_^ there, I hope it's right

OpenStudy (anonymous):

as simple as that :)

OpenStudy (anonymous):

sstar is completely right though i have a question where's the person who ASKED this question

OpenStudy (anonymous):

LOL! watching us, hey oneprince did you understand what we did? ^_^

OpenStudy (anonymous):

wait...he left ._.

OpenStudy (anonymous):

yeah -_- but you got it sstar i have integrating cosine sins T-T so messy...

OpenStudy (anonymous):

\[I=\frac{1}{3}(\frac{1}{32} (12 u-8 \sin(2 u)+\sin(4 u))+c)\]where u = 3x.

OpenStudy (anonymous):

o_o....no wait O_O!

OpenStudy (anonymous):

that's the final answer?

OpenStudy (anonymous):

Sub. u=3x, then du = 3dx --> dx = du/3. Then you have 1/3*int[(sin^4(u) du]

OpenStudy (anonymous):

You can use a reduction formula on sin^4(u)...

OpenStudy (anonymous):

no need for reduction formula ^_^ + i got you lol :)

OpenStudy (amistre64):

what happens when we derive:\[- \frac{\cos^5 (3x)}{15} + C\] ??

OpenStudy (anonymous):

why do you want to derive it? that's not the answer lol? ._.

OpenStudy (anonymous):

He wants to see if it matches the integrand.

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

lol.... if its not the answer, then what does it derive to?

OpenStudy (anonymous):

there's only one way to find out, T.R.Y ^_^

OpenStudy (anonymous):

give me a sec, i'll do it

OpenStudy (anonymous):

I won't even try since my page is lagging BADLY

OpenStudy (anonymous):

anyhow, I'm off to review for my DM test tomorrow, good luck all ^_^

OpenStudy (anonymous):

225 cos(3x)*sin(3x)^4

OpenStudy (anonymous):

Omg, I gave you the answer ^^

OpenStudy (anonymous):

wait i lied! i did * 15 instead of / 15 give me another min :P

OpenStudy (anonymous):

what answer ._.

OpenStudy (anonymous):

ohh i see why you put that 15 there it's cos(3x)*sin(3x)^4

OpenStudy (anonymous):

cheers :D

OpenStudy (anonymous):

Aren't you trying to find the integral of sin^4(3x)?

OpenStudy (anonymous):

yeah but amistre is trying to show that he's right but i showed that he's wrong ^^

OpenStudy (amistre64):

lol...not trying to show that im right, just trying to figure out what I did wrong :)

OpenStudy (anonymous):

:x sorry i didn't mean it like that but do you get it now?

OpenStudy (amistre64):

I see where my error cropped up ;)

OpenStudy (anonymous):

lols kewls